559 m 5. A student standing on the building at height 50.0 meter, shooting a yellow ball with velocity 20 m/s up, another student standing with him together, at same time, shooting a blue ball with velocity -20 m/s down. When does the yellow ball hit the ground? When does the blue ball hit the ground? What is yellow ball touch down velocity? What is blue ball touch down velocity?
Answer #1

Using second equation of motion,

s = ut + 0.5at^2

for yellow ball, -50 = 20*t + 0.5*-9.8*t^2

or t = 5.831 s

Yellow ball hits ground after 5.831 s

   for blue ball, -50 = -20*t + 0.5*-9.8*t^2

or t = 1.75 s

Blue ball hits ground after 1.75 s

Final velocity of yellow ball = 20+5.831*-9.8 = -37.15 m/s

Final velocity of blue ball = -20+1.75*-9.8 = -37.15 m/s